The main feature of the seismic activity in the period under review was the continuation of the low to moderate broad band tremor that commenced at 03:00 on 13 February. There was a slight increase in amplitude observed about 18:30 on 13 February and again at 03:00 on 14 February. This increased amplitude remained for about 1 hour and then began to slowly decrease in amplitude reaching its previous moderate level and has stayed so until the end of the period. Larger hybrid earthquakes (approximate magnitude 1.0) were observed every few hours on the nearest seismic stations to the volcano (Gages and Chances Peak) but were still too small to constrain instrumental locations. Other seismic signals often associated with small rockfalls and eruptions were observed throughout the period. Slightly larger signals of this type were observed at 22:56 on 13 February and 10:32 on 14 February.
EDM measurements were made on the eastern triangle to Castle Peak today. There was no change in slant distances on this triangle since its last occupation on 12 February.
Brief views of the crater were obtained from the helicopter late this afternoon. Areas of dome growth within English's Crater remains the same in the south-east and western parts of the dome. Several small rockfalls and minor steam emissions were observed from the south-east part of the dome. Steam venting continues to be mainly from the top central parts of the dome. Poor visibility due to low cloud did not permit clear views of the western part of the dome.
Dr Rick Hoblitt of the USGS and Mr Wilkie Balgobin of the SRU returned to their respective home bases today. Dr Hoblitt has been in Montserrat assisting with monitoring and interpretation of volcanic activity since early January. Mr Balgobin arrived on 26 January and assisted with the establishment and maintenance of seismic stations.